OpenStudy (anonymous):

what is the limit of the series as n approaches infinity of a(n) = (-3)^n / n! ? i know if it was +3 the limit approaches 0 using the squeeze theorem

terenzreignz (terenzreignz):

You can squeeze this too, you know ;) \[\huge \frac{-3^n}{n!}\le\frac{(-3)^n}{n!}\le\frac{3^n}{n!}\]

OpenStudy (anonymous):

so whats the final answer then?

terenzreignz (terenzreignz):

You tell me~ It's always in the middle of two sequences... BOTH of which go to zero. What do you think? :)

OpenStudy (anonymous):

how did u evaluate that -3^n/n! is equal to 0

terenzreignz (terenzreignz):

Because \[\huge \frac{3^n}{n!}\rightarrow 0\]Then if you take its negative, it would go to the negative of the limit, but then again, -0 = 0

OpenStudy (anonymous):

oh ok so i did not know u could take the negative of the limit ;)

OpenStudy (anonymous):

what if u made the left hand inequality just 0, would it be valid

terenzreignz (terenzreignz):

No, because sometimes, your sequence is less than zero. You need a sequence that is less than that at every turn.

OpenStudy (anonymous):

oh ok now i understand, thanks a lot man :)

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

just 1 more query before i close this question, can u please post your method of evaluating the limit of 3^n/n! is equal to 0. I want to see if it is simpler than the one taught to me

terenzreignz (terenzreignz):

Sorry, i was occupied @babycry hang on...

terenzreignz (terenzreignz):

Do you mean an epsilon proof?

terenzreignz (terenzreignz):

Coz I don't have one :)

OpenStudy (anonymous):

ok its ok then :)